Solve the system.
x + 2y = 10
y = 2x

Answer by checkley71(8403)

y=2x now substitute (2x) for y in the other equation & solve for x:
x+2(2x)=10
x+4x=10
5x=10
x=10/5
x=2 answer. now replace x by 2 & solve for y:
y=2*2
y=4 answer.
proof
2+2*4=10
2+8=10
10=10